OpenStudy (anonymous):

Side a=12 measure of angle B = 70 measure of angle C= 15 How do I find the answer using Law of Sines?

OpenStudy (precal):

for the laws of sine you need one angle and side. 180 - 70-15=95 degrees is angle a a/sin(A) = b/sin(B) small letters denote sides, Capital letters denote Angles

OpenStudy (precal):

To find side c use a/sin(A)=c/sin(C)

OpenStudy (anonymous):

Thanks! Your a life saver!

OpenStudy (precal):

Make sure your calculator is in degree mode.

OpenStudy (anonymous):

Always, if not my teacher would kill me!

OpenStudy (precal):

Well in precal, we go back and forth from degrees to radians, In calculus, we stay in radian.

OpenStudy (anonymous):

Wow in geometry we just stay in degree because we only use the calc of Sine, Cosine, and Tangent.

OpenStudy (precal):

Yes, I forgot in geometry they also, stay in the land of degrees. If you plan to study the higher mathematics, you will have to visit the land of radian.